OpenStudy (anonymous):

find the slope of the line tangent to the graph f(X) =5x^2 + 3x - 4 at x = 4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 5x^2 + 3x - 4\] \[\Large f(x+h) = 5(x+h)^2 + 3(x+h) - 4\] \[\Large f(x+h) = 5(x^2+2xh+h^2) + 3(x+h) - 4\] \[\Large f(x+h) = 5x^2+10xh+5h^2 + 3x+3h - 4\]

jimthompson5910 (jim_thompson5910):

\[\Large \lim_{h\to0}\frac{f(x+h)-f(x)}{h}\] \[\Large \lim_{h\to0}\frac{5x^2+10xh+5h^2 + 3x+3h - 4-(5x^2 + 3x - 4)}{h}\] \[\Large \lim_{h\to0}\frac{5x^2+10xh+5h^2 + 3x+3h - 4-5x^2 - 3x + 4}{h}\] \[\Large \lim_{h\to0}\frac{10xh+5h^2+3h}{h}\]

jimthompson5910 (jim_thompson5910):

\[\Large \lim_{h\to0}\frac{10xh+5h^2+3h}{h}\] \[\Large \lim_{h\to0}\frac{h(10x+5h+3)}{h}\] \[\Large \lim_{h\to0}10x+5h+3\] \[\Large 10x+5(0)+3\] \[\Large 10x+3\]

OpenStudy (anonymous):

find the derivative algebraically

OpenStudy (anonymous):

g(x) = 1/x at x = 2

OpenStudy (anonymous):

i found the equation of the tangent line. it is y = -1/4 (x-2) + 1/2 but idk how to find the derivate form the tangent line

jimthompson5910 (jim_thompson5910):

how did you find the equation of the tangent line?

OpenStudy (anonymous):

plugged it into y -y1 = m (x-x1)

OpenStudy (anonymous):

its correct right?

jimthompson5910 (jim_thompson5910):

how did you know the slope of the tangent line?

OpenStudy (anonymous):

i found it by doing the limit

jimthompson5910 (jim_thompson5910):

ok let me verify

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

ok I got \[\Large g^{\prime}(x)=-\frac{1}{x^2}\] So \[\Large g^{\prime}(2)=-\frac{1}{2^2}\] \[\Large g^{\prime}(2)=-\frac{1}{4}\] which means that the slope of the tangent line at x = 2 is -1/4 So you are correct there

OpenStudy (anonymous):

so what is the derivative

jimthompson5910 (jim_thompson5910):

g'(x) is the derivative function so in general, the derivative of g(x) is \[\Large g^{\prime}(x)=-\frac{1}{x^2}\] the derivative at x = 2 is the slope of the tangent line at x = 2 (so you just plug in x = 2 into g'(x))

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

how do you reduce this

OpenStudy (anonymous):

\[(1/(x+h)^{2} - 1/x ^{2}) / h\]

OpenStudy (anonymous):

\[3 = (102-100)/(35-y)\]